OpenStudy (anonymous):

is /36 a ratinol or irratinol #?

OpenStudy (anonymous):

is it a fraction? yes. then it is rational..... ANY FRACTION MEANS IT IS RATIONAL.... after at least 5 of these i assumed it would be evident.

OpenStudy (anonymous):

that is not a fraction

OpenStudy (anonymous):

is it \[\sqrt{36}=\sqrt{6}*\sqrt{6}=6\]? In this case it is rational

OpenStudy (anonymous):

thanks zed

OpenStudy (anonymous):

...................... it is /36. SOMETHING is being divided by 36 therefore it is a fraction. that is not a square root sign zed.

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

@RONNCC that's why i asked in case it was a typo. :) bluestar do you mean \[\frac{1}{36}\]or\[\sqrt{36}\]?, it is a little unclear